A bicyclist rode into the country for 5 h. In returning, her
speed was 5 mi/h faster and the trip took 4 h. What was her speed each way?

Answer by ankor@dixie-net.com(12692)

A bicyclist rode into the country for 5 h. In returning, her
speed was 5 mi/h faster and the trip took 4 h. What was her speed each way?
:
Let s = initial speed, (s+5) = return speed
:
We know the trip there, and the trip back are equal distances:
dist = time * speed
:
Dist out = Dist back
5s = 4(s+5)
5s = 4s + 20
5s - 4s = 20
s = 20 mph is her intital speed:
Obviously, 25 mph is her return speed:
:
Check using dist:
5(20) = 4(25)
:
Make sense to you?